OpenStudy (anonymous):

sin inverse(sine 9pi/7) Trying to find the exact value?

OpenStudy (anonymous):

It is just 9pi/7. Because sin^(-1) ( (sin x)) = x

OpenStudy (anonymous):

Even though 9pi/7 is bigger than pi/2?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Becuase you're back where you started from. Example: sin^(-1) (sin 30) = ? well, sin 30 = 1/2 and, sin^-1 (1/2) = 30..so we're back where we started from.

OpenStudy (anonymous):

The sin 9pi/7 on my calculator is -.7819?

OpenStudy (anonymous):

I'm confused on how to I know which is the answer... wouldn't it be -2pi/7 then since its the inverse of the sin9pi/7?

OpenStudy (anonymous):

Keep in mind, that sin^(-1) (x) is restricted between -pi/2 and pi/2 inclusive. Thats the reason for the discrepancy.

OpenStudy (anonymous):

The problem says... Find the exact value, in radians, of the expression of inverse sin(sin (9pi/7))

OpenStudy (anonymous):

? its the sin inverse of (sin(9pi/7))

OpenStudy (anonymous):

Since sin (x) and sin^(-1) (x) are inverse functions, then by definition, sin (sin^(-1)(x)) = x and (sin^(-1))(sin x) = x

OpenStudy (anonymous):

Easyaspi so then it would be the second definition? and the answer is simply 9pi/7? it seems to easy?

OpenStudy (anonymous):

dhord; What is critical here, is the restriction of Arcsin (x) or sin^(-1)(x). And that is it MUST be between -pi/2 and pi/2, inclusive. Thats why there's no problem with sin^(-1)(sin 30) = 30; and sin(sin^(-1)[1/2) = 1/2..because it falls between -pi/2 and pi/2. Look at the graph of y = sin^(-1)(x). It clearly shows that. But in 9pi/7, thats out of the restricted area, and thats where the problem occurs.

OpenStudy (anonymous):

\[\sin^{-1} (\sin( 9\pi/ 7 )\]

OpenStudy (anonymous):

that's the exact expression its asking to find the exact value in radians

OpenStudy (anonymous):

Is there no solution then? since its outside of pi/2

OpenStudy (mertsj):

\[\sin ^{-1}(\sin \frac{9\pi}{7})=\frac{-2\pi}{7}\]

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

If I may interject, sin (9pi/7) = - sin (2pi/7) by definition of third quadrant angle for siine. That is, 9pi/7, is a third-quadrant angle, so its value is computed by subtracting pi radians, you get 2pi/7, but since sine is negative in third quadrant, its value is -2pi/7.

OpenStudy (anonymous):

oh ok that makes more since

OpenStudy (anonymous):

So now you can find sin^(-1) becuase the argument is -2pi/7, which is within the restriction of sin^(-1)(x).

OpenStudy (anonymous):

I think the restriction of pi/2 was confusing me but now it makes since

OpenStudy (anonymous):

sense*

OpenStudy (anonymous):

:0